OpenStudy (anonymous):

Write an equation of a line that passes through the point (7, 3) and is parallel to the line y = negative 2 over 3x + 3. Y=3/2x-3 Y=3/2x+3 Y= -2/3x+23/3 Y= -2/3x-23/3

OpenStudy (misssunshinexxoxo):

I believe it is B

OpenStudy (phi):

parallel means they have the same slope Do you know what the slope of the second line is ?

OpenStudy (phi):

If you see an equation that matches the pattern y = mx + b where m is a number or fraction, and b is a number then you know m is the *slope*

OpenStudy (phi):

example y = 3x +2 slope is 3 (notice 3 is the number next to the "x") y= (1/3)x +17 slope is 1/3 (notice 1/3 is the number next to the x)

OpenStudy (phi):

your line is y= negative 2 over 3x + 3. it is hard to see the pattern when you use words. Write it with numbers y = -2/3 x + 3 what is the slope ?